### v2

# NeuralOM: Neural Ocean Model for Subseasonal-to-Seasonal Simulation

Abstract
Accurate Subseasonal-to-Seasonal (S2S) ocean simulation is critically important for marine research, yet remains challenging due to its substantial thermal inertia and extended time delay. Machine learning (ML)-based models have demonstrated significant advancements in simulation accuracy and computational efficiency compared to traditional numerical methods. Nevertheless, a significant limitation of current ML models for S2S ocean simulation is their inadequate incorporation of physical consistency and the slow-changing properties of the ocean system. In this work, we propose a neural ocean model (NeuralOM) for S2S ocean simulation with a multi-scale interactive graph neural network to emulate diverse physical phenomena associated with ocean systems effectively. Specifically, we propose a multi-stage framework tailored to model the ocean’s slowly changing nature. Additionally, we introduce a multi-scale interactive messaging module to capture complex dynamical behaviors, such as gradient changes and multiplicative coupling relationships inherent in ocean dynamics. Extensive experimental evaluations confirm that our proposed NeuralOM outperforms state-of-the-art models in S2S and extreme event simulation. The codes are available at https://github.com/YuanGao-YG/NeuralOM.
1 Introduction
Subseasonal-to-Seasonal (S2S) simulation represents a simulation of a dynamic system 2 to 6 weeks in advance liu2025cirt . Accurate S2S ocean simulation plays a pivotal role in marine research and activities, yet it remains a significant challenge due to the ocean’s substantial thermal inertia and prolonged time delay subramanian2019ocean ; lee2025observed . Traditional numerical ocean modeling systems, such as the Mercator Ocean Physical System (PSY4) and Real-Time Ocean Forecast System (RTOFS), rely on physics-driven numerical models to simulate ocean phenomena. These models use supercomputers to solve complex partial differential equations based on current ocean conditions francis2020high . Although these numerical systems achieve high accuracy, they are computationally expensive and time-consuming, requiring hours of computation even with hundreds of computational nodes. Moreover, enhancing the accuracy of these numerical methods is highly challenging, as they depend significantly on human expertise in understanding the complex physical laws governing the ocean environment.
Machine Learning (ML) provides a new perspective for ocean simulation due to several orders of magnitude faster than traditional numerical methods and high efficiency liu2024foundation  hao2025deep . Some spatio-temporal prediction algorithms, such as NMO wu2024neural , neXtSIM finn2024generative , OceanVP shi2024oceanvp focus on regional ocean modeling. And some ML-based task-specific ocean models are proposed, such as ENSO (El Niño/Southern Oscillation) prediction ham2019deep  chen2025combined , MJO (Madden Julian Oscillation) prediction kim2021deep  delaunay2022interpretable  shin2024deep , and MHWs (Marine Heatwaves) prediction shi2024physics  jacox2022global  shu2025advanced . In recent years, learning from the algorithms of ML-based global atmosphere foundation models pathak2022FourCastNet  bi2023accurate , ML-based global ocean foundation models have emerged, such as AI-GOMS xiong2023ai , Xihe wang2024xihe , and WenHai cui2025forecasting . Some foundation models achieve better performance than traditional numerical models. However, ML-based global ocean foundation models still face several core challenges: ❶ Slow-changing processes are hard to model. Due to substantial thermal inertia and extended time delay of the ocean system, it is difficult to learn the small changes effectively. ❷ Subseasonal-to-seasonal simulation lacks of adaptive modeling. Different time scales incorporate multi-scale ocean movements. Focusing on single-scale movement is hard to model consistently with the evolution process of ocean dynamics.
To address these challenges, this work proposes NeuralOM, a neural ocean model for subseasonal-to-seasonal simulation. Specially, to model the ocean’s slowly changing nature, we propose an ocean-specific multi-stage framework to learn large-scale to small-scale characteristics step by step. Under this framework, variables with seasonal periodicity, such as sea salinity, sea temperature, and sea surface height, are first subtracted from the climatology so that the model can better capture subtle changes. Additionally, we introduce a multi-scale interactive messaging module to capture complex dynamical behaviors, such as gradient changes and multiplicative coupling relationships inherent in ocean dynamics, which contributes to the S2S time scale simulation by adaptively modeling multi-scale movements.
The contribution of this work can be summarized as: (1) Stable Subseasonal-to-Seasonal Ocean Simulation Framework. We proposed an ocean-specific multi-stage framework tailored for the slow change nature of the ocean dynamic system, which achieves stale ocean simulation at S2S time scale. (2) Adaptively Modeling of Multi-scale Movements for Ocean Dynamic System. We propose a multi-scale interactive graph neural network to adaptively model the evolution process of ocean dynamic system. (3) Competitive Experimental Results. Extensive experiments demonstrate that our proposed model outperforms state-of-the-art models in S2S and extreme event simulation.
2 Related Work
ML-based Task-specific Ocean Model. The goal for this type of ocean model is to simulate specific ocean phenomena. Recent achievements includes ENSO (El Niño/Southern Oscillation) pediction ham2019deep  hu2021deep  chang2023predicting  lyu2024resonet  chen2025combined , MJO (Madden Julian Oscillation) prediction kim2021deep  delaunay2022interpretable  yang2023exploring  shin2024data  shin2024deep , MHWs (Marine Heatwaves) prediction jacox2022global  sun2024deep  parasyris2025marine  shu2025advanced , Sea Surface Height anomaly (SSHa) prediction song2020deep  liu2020sea . These models have demonstrated promising performance in comparison to traditional methods. However, several models overlook the interactions between multiple variables, suggesting that ML-based ocean models could potentially achieve enhanced accuracy by incorporating multi-variable interactions across 3D depth levels with a foundation model.
ML-based Global Ocean Foundation Model. The global ocean foundation models have significant achievements with higher accuracy and better efficiency than traditional numerical models. AI-GOMS xiong2023ai is an ocean foundation model that employs MAE he2022masked to simulate a multi-layer ocean state. Xihe wang2024xihe achieves an eddy-resolving simulation based on the hierarchical transformer. Some research, such as ORCA-DL guo2024data , OCE subel2024building , and DLESyM cresswell2024deep , targets decadal to climate-scale simulation, which indicates tremendous potential of ML-based models in climate simulation. LangYa yang2024langya is a cross-spatiotemporal ocean model with large-language model-based time embedding. Subsequently, WenHai cui2025forecasting incorporates domain knowledge into Swin Transformer liu2021swin and achieves better performance than the numerical model.
3 Method
Problem Setup In this work, we formulate ocean simulation as an autoregressive problem. For each time step $t$, we use the ocean state and forcing to simulate the next time step ocean state. The ocean state is composed of ocean depth level variables $\mathbf{O}^{d}_{t}$ and surface variables $\mathbf{O}^{s}_{t}$, while the forcing includes atmosphere surface variables $\mathbf{F}_{t}$ and $\mathbf{F}_{t+1}$. These components are concatenated along the channel dimension to form the combined input: $\mathbf{X}_{t}=[\mathbf{O}^{d}_{t},\mathbf{O}^{s}_{t},\mathbf{F}_{t},\mathbf{F}_{t+1}]\in\mathbb{R}^{N\times d}$, where $N=H\times W$ represents the number of grid locations, and $d=d_{d}+d_{s}+d_{f}$ is the total number of input variables. Here, $d_{d}$, $d_{s}$, and $d_{f}$ are the numbers of ocean depth level variables, surface level variables, and forcing variables, respectively. In our setup, the input contains 101 variables: 92 depth level variables, 1 surface level variable, and 8 forcing variables. The model aims to simulate the next time ocean state $\hat{\mathbf{O}}_{t+1}$, given the current input $\mathbf{X}_{t}$, capturing the spatiotemporal evolution of the ocean: $\hat{\mathbf{O}}_{t+1}=\text{Model}(\mathbf{X}_{t};\Theta),$ where $\Theta$ denotes the model parameters. $\hat{\mathbf{O}}_{t+1}=[\hat{\mathbf{O}}^{d}_{t+1},\hat{\mathbf{O}}^{s}_{t+1}]\in\mathbb{R}^{N\times d_{o}}$, where $d_{o}=d_{d}+d_{s}$ is the total number of ocean variables. The training objective is to minimize the relative mean squared error (MSE) between the simulations and the true values across all time steps: $\min_{\Theta}\frac{1}{T}\sum_{t=0}^{T-1}\frac{\left\|\hat{\mathbf{O}}_{t+1}-\mathbf{O}_{t+1}\right\|_{2}^{2}}{\left\|\mathbf{O}_{t+1}\right\|_{2}^{2}}$. During inference, we employ a rollout strategy to simulate longer sequences. Starting from the initial state $\mathbf{X}_{0}$, the model recursively uses its previous simulations as the next input: $\hat{\mathbf{O}}_{t+1}=\text{Model}(\hat{\mathbf{X}}_{t};\Theta),\quad t=0,1,2,\dots,T-1.$
3.1 Ocean-specific Multi-stage Framework
Due to its substantial thermal inertia and extended time delay, the ocean changes slowly compared to its original state, making it difficult to effectively learn and capture changes. Therefore, as depicted in Figure 1(a), we propose an ocean-specific multi-stage framework (OMF) to learn small change nature. Specifically, different from previous works such as AI-GOMS, Xihe, and WenHai, we first remove the climatological mean from variables exhibiting strong periodicity before feeding them into the network, thereby directing the model’s attention toward small anomalies:
$O_{t}^{periodic}={O_{t}}^{original}-{O_{t}}^{climatology},$ (1)
where, $O_{t}^{periodic}$ denotes periodic ocean variables (sea salinity, sea temperature, and sea surface height), ${O_{t}}^{original}$ represents the original ocean variables, ${O_{t}}^{climatology}$ is the climatology calculated by the historical data from 1993 to 2017, exclude from valid and test data. Therefore, the input ocean variables ${O}_{t}=[{O}^{periodic}_{t},{O}^{u,v}_{t}]\in\mathbb{R}^{N\times d_{o}}$. ${O}^{u,v}_{t}$ is sea stream velocity. Then, inspired by mojgani2024interpretable that intends to reduce step error for scientific data prediction and a multi-step supervised method trask2022hierarchical , we adopt a stage-wise learning strategy to sequentially capture the ocean’s gradual evolution. It is worth noting that although the previous works could reduce single-step error for some ideal datasets, they are hard to achieve long-term simulation in complex real-world ocean scenes. In contrast, our proposed OMF focuses on long-term simulation and achieves stable subseasonal-to-seasonal ocean simulation. Specifically, the intermediate results of the next step ocean state $\hat{O}_{t+1}$ can be acquired using the combined input $X_{t}=[O_{t},F_{t},F_{t+1}]\in\mathbb{R}^{N\times d}$:
$\hat{O}^{1}_{t+1}={\mathcal{F}^{M}_{1}}(X_{t}),$ (2)
where, ${\mathcal{F}^{M}_{1}}$ is the first model with M-step finetune, which can be trained step by step from the 1-step supervised model. Subsequently, as shown in Figure 1(a), we apply the second model to learn the residue between $\hat{O}^{1}_{t+1}$ and ${O}_{t+1}$, thereby getting the intermediate results produced by the second model:
$\hat{O}^{2}_{t+1}={\mathcal{F}^{N}_{2}}(\hat{O}^{1}_{t+1})+\hat{O}^{1}_{t+1},$ (3)
where, ${\mathcal{F}^{M}_{2}}$ is the second model with N-step finetune, which is trained step by step. For the last model:
$\hat{O}_{t+1}={\mathcal{F}^{N}_{Q}}(\hat{O}^{Q-1}_{t+1})+\hat{O}^{Q-1}_{t+1},$ (4)
where, ${\mathcal{F}^{N}_{Q}}$ denotes the Qth model with N-step finetune. In our experiment, we set M=6, N=10, and Q=2 as an example to demonstrate that the proposed ocean-specific multi-stage framework could improve the long-term simulation for complex ocean dynamic systems.
3.2 Multi-scale Interactive Graph Neural Networks
3.2.1 Graph Encoder
Following the previous works lam2023learning  gao2025oneforecast , we employ the multi-scale graph encoder to map the data from lat-lon grids into a graph structure. The multi-scale graph structure can be defined as:
$\mathcal{G}=\left(\mathcal{V}^{G},\mathcal{V},\mathcal{E}_{m},\mathcal{E}^{\mathrm{G}2\mathrm{M}},\mathcal{E}^{\mathrm{M}2\mathrm{G}}\right),$ (5)
where, ${\mathcal{V}}^{G}$ represents the set of lat-lon grid nodes, with a total of $N=H\times W$ nodes; $\mathcal{V}$ represents the mesh nodes; $\mathcal{E}_{m}$ denotes the multi-scale edges, which represents the edge with different lengths; $\mathcal{E}^{\mathrm{G}2\mathrm{M}}$ and $\mathcal{E}^{\mathrm{M}2\mathrm{G}}$ are the unidirectional edges that connect lat-lon grid nodes and mesh nodes. All scales share the same set of nodes $\mathcal{V}$ but with multi-level edges. Then, we apply an MLP to map the data to the latent space, which can be defined as:
$(\mathcal{V}_{f}^{G},h,\mathcal{E}_{f},\mathcal{E}^{\mathrm{G}2\mathrm{M}}_{f},\mathcal{E}^{\mathrm{M}2\mathrm{G}}_{f})=\mathrm{MLP}(X_{t},\mathcal{V},\mathcal{E}_{m},\mathcal{E}^{\mathrm{G}2\mathrm{M}},\mathcal{E}^{\mathrm{M}2\mathrm{G}}),$ (6)
$MLP=\mathcal{N}(\sigma(\mathcal{L}(\cdot))),$ (7)
where, $\mathcal{L(\cdot)}$ is the linear function. $\sigma(\cdot)$ is the SiLU activation function, $\mathcal{N}(\cdot)$ is the LayerNorm function. $X_{t}$, $\mathcal{V}$, $\mathcal{E}$, $\mathcal{E}^{\mathrm{G}2\mathrm{M}}$, and $\mathcal{E}^{\mathrm{M}2\mathrm{G}}$ are embedded features of grid nodes, mesh nodes, mesh edges, grid to mesh edges, and mesh to grid edges. Then, we update the grid2mesh edge features using information from the adjacent nodes:
${\mathcal{E}^{\mathrm{G}2\mathrm{M}}_{f}}^{\prime}=\mathrm{ESMLP}(\mathcal{V}_{f}^{G},h,\mathcal{E}^{\mathrm{G}2\mathrm{M}}_{f}),$ (8)
where, $\mathrm{ESMLP}(\cdot)$ is the Edge Sum MLP pfaff2020learning . The mesh node features are updated by an MLP:
$h^{\prime}=\mathrm{MLP_{e1}}(h,\sum{\mathcal{E}^{\mathrm{G}2\mathrm{M}}_{f}}^{\prime}),$ (9)
where, $\sum\mathcal{E}_{f}^{\mathrm{G}2\mathrm{M}^{\prime}}$ are the edges that arrives at mesh node. The grid node features are updated using another MLP:
$\mathcal{V}_{f}^{G^{\prime}}=\operatorname{MLP}_{e2}\left(\mathcal{V}_{f}^{G}\right).$ (10)
Finally, the residual connections are applied to update $\mathcal{E}^{\mathrm{G}2\mathrm{M}}_{f}$, $h$, and $\mathcal{V}_{f}^{G}$ again:
${\mathcal{E}^{\mathrm{G}2\mathrm{M}}_{f}}={\mathcal{E}^{\mathrm{G}2\mathrm{M}}_{f}}+{\mathcal{E}^{\mathrm{G}2\mathrm{M}}_{f}}^{\prime},h=h+h^{\prime},\mathcal{V}_{f}^{G}=\mathcal{V}_{f}^{G}+{\mathcal{V}_{f}^{G}}^{\prime}.$ (11)
Figure 1: Overview of Our NeuralOM. (a) The overall architecture of the ocean-specific multi-stage framework, input variables (subtracting climatology for periodic variables), multi-model stage, and visualization of simulation ocean variables; (b) The proposed multi-scale interactive graph neural networks; (c) The global ocean simulation module uses a rollout approach to generate future results.
3.2.2 Multi-scale Interactive Massaging
To make the message passing more consistent with the evolution process of the ocean dynamics system, we propose a module called Multi-scale Interactive Messaging (MIM). MIM includes a physical process-based edge update module and a multi-scale node update module. In our setup, 16 MIMs are applied to conduct massaging.
Physical Process-based Edge Update Module. In ocean simulation, ocean currents, temperature, salinity, and other factors interact with each other, and these interactions span different scales and layers. Unlike traditional MLP-based update strategy, which update edges by simply concatenating edges and nodes, we introduce a physical process-based edge update module to more accurately simulate the dynamic relationships in the ocean system. For the sender node feature $\mathbf{h}_{s(i)}$ and receiver nodes feature $\mathbf{h}_{r(i)}$, we first calculate their mutual interaction:
$\mathbf{h}_{d(i)}=\mathbf{h}_{s(i)}-\mathbf{h}_{r(i)},$ (12)
where, $\mathbf{h}_{d(i)}$ represents the variation in features between neighboring nodes, capturing the gradient changes of quantities such as flow velocity and salinity in the ocean. For example, when simulating changes in ocean currents, the velocity differences can indicate the energy transfer and dynamic characteristics of different ocean regions.
$\mathbf{h}_{mp(i)}=\mathbf{h}_{s(i)}*\mathbf{h}_{r(i)},$ (13)
where, $\mathbf{h}_{mp(i)}$ represents the multiplicative coupling relationship between nodes, similar to the coupling phenomenon between temperature and salinity in the ocean. The temperature and salinity of seawater together determine its density, which is crucial for modeling deep ocean currents.
$\mathbf{h}_{cos(i)}=cos(\mathbf{h}_{s(i)},\mathbf{h}_{r(i)}),$ (14)
where, $\mathbf{h}_{cos(i)}$ is the cosine similarity between different nodes. For the ocean systems, it can be used to identify similar water masses or marine subsystems. For instance, when the changes in seawater temperature and salinity become consistent, it may indicate that these two ocean regions are influenced by similar climate systems, allowing information to be transmitted through similarity and improving the accuracy of the simulation. We define a function $\mathcal{R}(\cdot)$ to fuse mesh node features and the features produced by the interaction:
$\mathcal{R}(x)=\mathcal{\sigma}(\mathcal{N}(\mathcal{L}(\mathcal{C}(x)))),$ (15)
where, $\mathcal{C}$ represents the concatenate function. Then, the fused mesh node features can be obtained by:
$\mathbf{h}_{s(i)}=\mathcal{R}(\mathbf{h}_{s(i)},\mathbf{h}_{mp(i)},\mathbf{h}_{mp(i)},\mathbf{h}_{cos(i)}),\mathbf{h}_{r(i)}=\mathcal{R}(\mathbf{h}_{r(i)},\mathbf{h}_{mp(i)},\mathbf{h}_{mp(i)},\mathbf{h}_{cos(i)}).$ (16)
Subsequently, the edge features are updated using the information from adjacent nodes:
$\mathcal{E}_{f}^{\prime}=\mathbf{W}_{e}\mathcal{E}_{f},h_{s}^{\prime}=\mathbf{W}_{s}h_{s},h_{r}^{\prime}=\mathbf{W}_{r}h_{r}+\mathbf{b}_{r},$ (17)
$\mathbf{h}_{\mathrm{sum}}=\mathcal{E}_{f}^{\prime}+h_{s}^{\prime}+h_{r}^{\prime},$ (18)
${\mathcal{E}_{f}}^{\prime}=\mathcal{N}\left(\mathbf{W}{\cdot}\sigma\left(h_{sum}\right)+\mathbf{b}\right),$ (19)
where, $\mathbf{W}_{e}$, $\mathbf{W}_{s}$, $\mathbf{W}_{r}$ are the linear transformation matrix of edge features, send node feature, and receive node features. $\mathbf{W}$ is the linear transformation matrix of output layer. $b_{r}$ and $b$ are the bias. Finally, the edge features are updated using the residual connection:
${\mathcal{E}_{f}}={\mathcal{E}_{f}}^{\prime}+{\mathcal{E}_{f}}.$ (20)
Multi-scale Node Update Module. Unlike traditional node update strategies, which typically sum or average the features of neighboring nodes, we consider the movements of the ocean at different scales. To achieve adaptive adjustment of aggregation methods across different scales, we introduce a multi-scale node update module that dynamically adjusts the weights of the aggregation strategy. For rapid local ocean dynamics with smaller scales, the gating mechanism automatically tends to favor sum aggregation, capturing the local rapid dynamic changes quickly and accurately. Conversely, when the model identifies that the motion scale around the node is larger and the changes are slower, it tends to prefer mean aggregation, which achieves a smoother and more robust representation of the large-scale ocean state.
Specifically, in the ocean system, meso- and small-scale movements (such as eddies, localized strong upwelling, coastal currents, etc.) exhibit distinct characteristics of localized energy concentration and rapid dynamic changes. In these scenarios, the influence of adjacent regions on a node is often cumulative, such as the concentrated propagation of local heat, momentum, or salt flux. In such cases, the sum aggregation strategy, by directly adding the contributions from surrounding nodes, more accurately reflects the rapid evolution and locally concentrated dynamic features, which can be expressed as:
$h_{sum}^{\prime}=\mathrm{MLP}_{sum}(h,\sum\mathcal{E}_{f}^{\prime}).$ (21)
For large-scale, slowly varying ocean movements (such as the global ocean circulation, thermohaline circulation, and long-term ocean current trends), the differences between nodes and the extent of local fluctuations are smaller, and information transfer between neighboring nodes tends to be in a relatively steady and averaged state. In this case, the mean aggregation strategy can more stably capture long-term trends and the overall balanced state, avoiding the over-influence of short-term local disturbances on the large-scale state, which can be expressed as:
$h_{mean}^{\prime}=\mathrm{MLP}_{mean}(h,\sum\mathcal{E}_{f}^{\prime}).$ (22)
Then, we use $h_{sum}^{\prime}$ and $h_{mean}^{\prime}$ to calculate gating coefficient:
$\gamma=\mathrm{MLP_{g}}(\mathcal{C}(h_{sum}^{\prime},h_{mean}^{\prime})).$ (23)
Finally, an MLP and a residual connection are used to update the node features:
$h^{\prime}=\mathrm{MLP_{node}}(\gamma{\cdot}h_{sum}^{\prime}+(1-\gamma){\cdot}h_{mean}^{\prime}).$ (24)
3.2.3 Decoder
The role of the decoder is to decode the information from the latent space to the lat-lon grid. We first use the information from the adjacent nodes to update mesh2grid features:
$\mathcal{E}_{f}^{\mathrm{M}2\mathrm{G}^{\prime}}=\operatorname{ESMLP}\left(\mathcal{V}_{f}^{G},h,\mathcal{E}_{f}^{\mathrm{M}2\mathrm{G}}\right).$ (25)
Then, we update the grid node features using the information of edges that arrive at the grid nodes:
${\mathcal{V}_{f}^{G}}^{\prime}={\mathrm{MLP}}_{d1}(\mathcal{V}_{f}^{G},\sum{\mathcal{E}^{\mathrm{M}2\mathrm{G}}_{f}}^{\prime}).$ (26)
Subsequently, we apply residual connection to update grid node features:
$\mathcal{V}_{f}^{G}=\mathcal{V}_{f}^{G}+\mathcal{V}_{f}^{G^{\prime}}.$ (27)
Finally, we use an MLP to predict the next step results:
$\hat{O}_{t+1}=\operatorname{MLP}_{d2}\left(\mathcal{V}_{f}^{G}\right).$ (28)
3.3 Optimization
We apply the relative $L_{2}$ loss function to train each component of NeuralOM, which can be defined as:
$\mathcal{L}_{2}=\frac{1}{KHW}\sum_{k=1}^{K}\sum_{i=1}^{H}\sum_{j=1}^{W}\frac{\left(\hat{O}_{i,j,k}^{t+l\delta t}-O_{i,j,k}^{t+l\delta t}\right)^{2}}{\left(O_{i,j,k}^{t+l\delta t}\right)^{2}},$ (29)
where, $\hat{O}_{i,j,k}^{t+l\delta t}$ and $O_{i,j,k}^{t+l\delta t}$ are the predicted and ground truth for ocean variables (channels) $k$ at spatial location $(i,j)$ and time $t+l\delta t$; $K$ is the number of variables; $H$ and $W$ are the height and width of the spatial dimensions, respectively; $\delta t$ is the time interval of single-step simulation (we use $\delta t=24$ hours).
3.4 Global Ocean Simulation
As illustrated in Figure 1(c), our inference procedure follows a rollout strategy: starting from the initial state $\mathbf{X}_{0}$, the trained model autoregressively feeds its own previous simulations together with the corresponding external forcing into the next time step. Previous studies have shown that evaluating an ocean model with forecast atmospheric fields can blur the distinctions among different ocean models, because errors in the atmospheric forecasts tend to dominate and “smooth out” their performance differences wang2024xihe  shu2025advanced . In other words, when an ocean-only model is driven by biased forecast forcing, a well-designed ocean model can appear worse, while a poorer one can appear better. Consequently, a more reliable assessment of an ocean model’s skill is obtained by using ground truth atmospheric forcing rather than forecast fields. Therefore, we use the ground truth atmospheric forcing to drive ocean models for comprehensive comparisons.
4 Experiment
In this section, we thoroughly evaluate the performance of NeuralOM, including metric results and visual results. We conduct all experiments on 64 NVIDIA A100 GPUs.
Table 1: In the global ocean simulation task, we compare the performance of our NeuralOM with 3 baselines. The average results for all 92 ocean variables of RMSE and ACC are recorded. Given the magnitude of the values for different variables, we report the normalized RMSE. A smaller RMSE ($\downarrow$) and a bigger ACC ($\uparrow$) indicate better performance. The best results are in bold, and the second best are with underline.
Model Metric
40-day 45-day 50-day 60-day
RMSE ACC RMSE ACC RMSE ACC RMSE ACC
FourCastNet pathak2022FourCastNet arXiv 2022 5.6336 0.0781 >10 0.0189 >10 0.0046 >10 0.0007
CirT liu2025cirt ICLR 2025 2.7525 0.0143 2.8700 0.0115 2.9528 0.0099 3.0645 0.0081
WenHai cui2025forecasting Nature Communications 2025 0.7200 0.5067 0.7468 0.4705 0.7702 0.4386 0.8093 0.3852
NeuralOM (ours) 0.6619 0.5210 0.6785 0.4916 0.6926 0.4656 0.7158 0.4212
NeuralOM (Promotion) 8.07% 2.84% 9.15% 4.49% 10.07% 6.16% 11.55% 9.33%
Figure 2: 60-day simulation results of different models.
Figure 3: The latitude-weighted RMSE (lower is better) and ACC (higher is better) results of several important ocean surface variables.
Table 2: Ablation Studies on whether subtracting the climatology from periodicity variables will improve the simulation performance, the best results are in bold.
VARIANTS 23 Layers S 23 Layers T SSH
RMSE ACC RMSE ACC RMSE ACC
NeuralOM w/o Subtract Climatology 0.0515 0.1061 0.1518 0.4535 0.2424 0.5747
NeuralOM W Subtract Climatology 0.0105 0.6412 0.0985 0.6508 0.1374 0.6695
4.1 Benchmarks and Baselines
We conduct the experiments on GlORYS12 reanalysis data, which offers daily mean data covering latitudes between -80° and 90° spanning between 1993 and 2023. The subset we use includes years from 1993 to 2020, which is 1993-2017 for training, 2018-2019 for validating, and 2020 for testing. We use 4 depth level ocean variables (each with 23 depth levels, corresponding to 0.49 m, 2.65 m, 5.08 m, 7.93 m, 11.41 m, 15.81 m, 21.60 m, 29.44 m, 40.34 m, 55.76 m, 77.85 m, 92.32 m, 109.73 m, 130.67 m, 155.85 m, 186.13 m, 222.48 m, 266.04 m, 318.13 m, 380.21 m, 453.94 m, 541.09 m and 643.57 m), Sea salintiy (S), Sea stream zonal velocity (U), Sea stream meridonal velocity (V), Sea temperature (T), and 1 surface level variable Sea surface height (SSH). We downsampling them to 1/2 degree (H=361, W=720). And we use the data with size 360 × 720. 4 atmosphere variables are used as forcing. More details can be found in the Appendix. We compare the simulation performance of NeuralOM with 3 data-driven models (FourCastNet pathak2022FourCastNet , CirT liu2025cirt , and WenHai cui2025forecasting ). FourCastNet is an atmospheric backbone model with good performance. CirT is a SOTA subseasonal-to-seasonal atmosphere prediction model that is similar to our simulation time scale. WenHai is a SOTA ML-based ocean model that outperforms traditional numerical models and ML-based models. Specifically, we directly use the inference code released by WenHai’s author to produce the simulation results of WenHai. For other baselines and our NeuralOM, we retrain them in the same framework and settings.
4.2 Evaluation Metrics
We utilize four metrics, RMSE (Root Mean Square Error), ACC (Anomalous Correlation Coefficient), CSI (Critical Success Index), and SEDI (Symmetric Extremal Dependence Index) to evaluate the simulation performance, which can be defined as:
$\operatorname{RMSE}(\mathcal{J},t)=\sqrt{\frac{\sum\limits_{i=1}^{N_{\text{lat}}}\sum\limits_{j=1}^{N_{\text{lon}}}L(i)\left(\hat{\mathbf{A}}_{ij,t}^{\mathcal{J}}-\mathbf{A}_{ij,t}^{\mathcal{J}}\right)^{2}}{N_{\text{lat}}\times N_{\text{lon}}}},$ (30)
$\operatorname{ACC}(\mathcal{J},t)=\frac{\sum\limits_{i=1}^{N_{\text{lat}}}\sum\limits_{j=1}^{N_{\text{lon}}}L(i)\hat{\mathbf{A^{\prime}}}_{ij,t}^{\mathcal{J}}\mathbf{A^{\prime}}_{ij,t}^{\mathcal{J}}}{\sqrt{\sum\limits_{i=1}^{N_{\text{lat}}}\sum\limits_{j=1}^{N_{\text{lon}}}L(i)\left(\hat{\mathbf{A^{\prime}}}_{ij,t}^{\mathcal{J}}\right)^{2}\times\sum\limits_{i=1}^{N_{\text{lat}}}\sum\limits_{j=1}^{N_{\text{lon}}}L(i)\left(\mathbf{A^{\prime}}_{ij,t}^{\mathcal{J}}\right)^{2}}},$ (31)
where $\mathbf{A}_{i,j,t}^{\mathcal{K}}$ represents the value of variable $\mathcal{J}$ at horizontal coordinate $(i,j)$ and time t. Latitude-dependent weights are defined as $L(i)=N_{\text{lat }}\times\frac{\cos\phi_{i}}{\sum_{i\textquoteright=1}^{N_{\text{lat}}}\cos\phi_{i\textquoteright}}$, where $\phi_{i}$ is the latitude at index i. The anomaly of $A$, denoted as $A^{\prime}$, is computed as the deviation from its climatology, which corresponds to the long-term mean of the meteorological state estimated from multiple years of training data. RMSE and ACC are averaged across all time steps and spatial coordinates, providing summary statistics for each variable $\mathcal{J}$ at a given lead time $\Delta t$.
$\operatorname{CSI}(\mathcal{J},t)=\frac{\mathrm{TP}}{\mathrm{TP}+\mathrm{FP}+\mathrm{FN}},$ (32)
$\operatorname{SEDI}(\mathcal{J},t)=\frac{\log(F)-\log(H)-\log(1-F)+\log(1-H)}{\log(F)+\log(H)+\log(1-F)+\log(1-H)},$ (33)
where, true positives (TP) represent the number of instances in which the ocean state is correctly simulated, while false positives (FP) and false negatives (FN) are defined analogously. $F=\frac{\mathrm{FP}}{\mathrm{FP}+\mathrm{TP}}$ is the false alarm rate, and $H=\frac{\mathrm{TP}}{\mathrm{TP}+\mathrm{FN}}$ represents the hit rate.
4.3 Global Ocean Simulation
Given the magnitude of the values for different variables, we report the normalized RMSE. We first normalize the 93 ocean variables and then calculate the indicators for the 240 initial conditions (ICs) starting from 1 Jan. 2020, and each IC will be updated one day later.
Figure 4: The CSI and SEDI results of extreme event assessment.
As shown in Table 1, NeuralOM achieves competitive performance compared with the state-of-the-art models. As shown in Figure 2, NeuralOM is closer to the ground truth. We also show the simulation results of several important surface variables in Figure 3, which are not normalized. It can be seen that state-of-the-art atmosphere prediction model, such as CirT, perform poorly in ocean scenes. A main reason may lays on the different nature of atmosphere and ocean dynamic systems. In contrast, our proposed NeuralOM performs satisfactorily in subseasonal-to-seasonal ocean simulation. This improvement is primarily attributed to integrating the proposed ocean-specific multi-stage framework and multi-scale interactive graph neural network, which effectively captures the ocean’s small change nature and accurately models the evolution of ocean dynamics at different time scales. More results at other depth-level layers can be found in the Appendix. It is worth noting that our NeuralOM achieves competitive results compared with the SOTA WenHai model training using 1/12 degree data, although we train NeuralOM using lower resolution data with 1/2 degree.
4.4 Extreme Events Assessment
We apply CSI and SEDI to evaluate the simulation ability for extreme events. As shown in Figure 4, we calculate the CSI and SEDI results of U0 and V0 for 50 ICs. A bigger value means better performance to extreme event simulation. Our proposed NeuralOM achieves competitive performance.
4.5 Ablation Studies
To assess the effectiveness of the proposed method, we conduct detailed ablation studies, as summarized in Table 2 and Table 3. Specially, we introduce several model variants: (1) NeuralOM W/O Subtract Climatology, we don’t subtract climatology for periodicity variables before input them to the model.
Table 3: Ablation Studies on model design, the best results are in bold.
Variants RMSE ACC
NeuralOM w/o OMF 0.6692 0.5503
NeuralOM w/o PEM 0.6865 0.5168
NeuralOM w/o MNM Sum 0.7037 0.5361
NeuralOM w/o MNM Mean 0.6978 0.5161
NeuralOM 0.6097 0.5983
(2) NeuralOM w Subtract Climatology, we subtract the climatology for the periodicity variables. (3) NeuralOM w/o OMF, we remove ocean-specific multi-stage framework. (4) NeuralOM w/o PEM, we remove physical process-based edge update module. (5) NeuralOM w/o MNM Sum, we remove the sum aggregation strategy in the multi-scale node update module. (6) NeuralOM w/o MNM Mean, we remove the mean aggregation strategy in the multi-scale node update module. (7) NeuralOM, the full version of NeuralOM. The above metrics are based on 30-day simulation results (50ICs for (1),(2), and 100ICs for others). Experimental results show that removing any single component impairs NeuralOM’s performance, thereby validating the effectiveness of our proposed method.
5 Conclusion
This work presents NeuralOM, a neural ocean foundation model tailored for Subseasonal-to-Seasonal (S2S) simulation. Our proposed ocean-specific multi-stage framework ensures accurate modeling of the ocean’s slow dynamics, while the multi-scale interactive messaging module enables the representation of intricate dynamical interactions, which accurately model the evolution process of ocean dynamics at the S2S time scale. Extensive experimental results demonstrate that NeuralOM significantly outperforms existing state-of-the-art models in both S2S and extreme event simulation, confirming its potential to revolutionize ocean and climate simulation tasks.
Acknowledgements
This work was supported by the National Natural Science Foundation of China (42125503, 42430602).
Appendix A Data Details
A.1 Dataset
We conduct the experiments on GlORYS12 reanalysis data, offering daily mean data covering latitudes between -80° and 90° spanning between 1993 and 2023, which can be downloaded from https://data.marine.copernicus.eu/product/GLOBAL_MULTIYEAR_PHY_001_030/description. The subset we use includes years from 1993 to 2020, which is 1993-2017 for training, 2018-2019 for validating, and 2020 for testing. We use 4 depth level ocean variables (each with 23 depth levels, corresponding to 0.49 m, 2.65 m, 5.08 m, 7.93 m, 11.41 m, 15.81 m, 21.60 m, 29.44 m, 40.34 m, 55.76 m, 77.85 m, 92.32 m, 109.73 m, 130.67 m, 155.85 m, 186.13 m, 222.48 m, 266.04 m, 318.13 m, 380.21 m, 453.94 m, 541.09 m and 643.57 m), Sea salinity (S), Sea stream zonal velocity (U), Sea stream meridional velocity (V), Sea temperature (T), and 1 surface level variable Sea surface height (SSH). To improve computational efficiency, we use bilinear interpolation to downsampling them to 1/2 degree (H=361, W=720). And to better adapt to the input of different architecture models, we use the data with size 360 × 720. 4 atmosphere variables from ERA5 are used as forcing, which include 10 metre u wind component (U10M), 10 metre v wind component (V10M), 2 metre temperature (T2M), and mean sea level pressure (MSLP). The ERA5 data can be downloaded from https://cds.climate.copernicus.eu/, the official website of Climate Data Store (CDS). All the data we used are shown in Table 4.
A.2 Data preprocessing
Different ocean and forcing variables exhibit substantial variations in magnitude. To enable the model to concentrate on accurate simulation rather than learning the inherent magnitude discrepancies among variables, we normalize the input data prior to model ingestion. Specifically, for ocean variables, we compute the mean and standard deviation from the training dataset spanning the years 1993 to 2017. For forcing variables, we calculate these statistics from an extended dataset covering the period 1959 to 2017. Each variable thus possesses a dedicated mean and standard deviation. Before inputting data into the model, we normalize the data by subtracting the corresponding mean and dividing the respective standard deviation. For the ‘nan’ values of land, we fill them with zero before inputting the data into the model.
Table 4: The data details in this work.
Type Full name Abbreviation Layers Time Dt Spatial Resolution
Atmosphere 10 metre u wind component U10M 1 1993-2020 24h 0.5°
Atmosphere 10 metre v wind component V10M 1 1993-2020 24h 0.5°
Atmosphere 2 metre temperature T2M 1 1993-2020 24h 0.5°
Atmosphere Mean sea level pressure MSLP 1 1993-2020 24h 0.5°
Ocean Sea salinity S 23 1993-2020 24h 0.5°
Ocean Sea stream zonal velocity U 23 1993-2020 24h 0.5°
Ocean Sea stream meridional velocity V 23 1993-2020 24h 0.5°
Ocean Sea temperature T 23 1993-2020 24h 0.5°
Ocean Sea surface height SSH 1 1993-2020 24h 0.5°
Static Land-sea mask LSM 0.5°
Appendix B Algorithm
We summarize the overall framework of the ocean-specific multi-stage framework (OMF) in Algorithm 1 and multi-scale interactive graph neural networks (MIGNN) in Algorithm 2.
Algorithm 1 Ocean-specific Multi-stage Framework (OMF)
0:  Initial input data $X_{t}$.
0:  Next step ocean state $O_{t+1}$.
1:  Initialize OMF
2:  repeat
3:     MIGNN (1)
4:     Apply the first MIGNN model to encode the combined input $X_{t}=[O_{t},F_{t},F_{t+1}]$ into the intermediate results of the next step ocean state $\hat{O}^{1}_{t+1}={\mathcal{F}^{M}_{1}}(X_{t})$.
5:     MIGNN (2)
6:     Apply the second MIGNN model to map $\hat{O}^{1}_{t+1}$ into the intermediate results: $\hat{O}^{2}_{t+1}={\mathcal{F}^{N}_{2}}(\hat{O}^{1}_{t+1})+\hat{O}^{1}_{t+1}$.
7:     MIGNN (Q)
8:     Apply the Q-th MIGNN model to map $\hat{O}^{Q-1}_{t+1}$ into the final results: ${O}_{t+1}={\mathcal{F}^{N}_{Q}}(\hat{O}^{Q-1}_{t+1})+\hat{O}^{Q-1}_{t+1}$.
9:  until converged
10:  return $OMF$
Algorithm 2 Multi-scale Interactive Graph Neural Networks (MIGNN)
0:  Initial input data $X_{t}$.
0:  Next step ocean state $O_{t+1}$.
1:  Initialize MIGNN
2:  repeat
3:     Encoder
4:     Embedding features of grid nodes $X_{t}$, mesh nodes $\mathcal{V}$, mesh edges $\mathcal{E}_{m}$, grid to mesh edges $\mathcal{E}^{\mathrm{G}2\mathrm{M}}$, and mesh to grid edges $\mathcal{E}^{\mathrm{M}2\mathrm{G}}$ into latent space using separate MLP: ($\mathcal{V}_{f}^{G}$, $h$, $\mathcal{E}_{f}$, $\mathcal{E}^{\mathrm{G}2\mathrm{M}}_{f}$, $\mathcal{E}^{\mathrm{M}2\mathrm{G}}_{f}$) = MLPs($X_{t}$, $\mathcal{V}$, $\mathcal{E}_{m}$, $\mathcal{E}^{\mathrm{G}2\mathrm{M}}$, $\mathcal{E}^{\mathrm{M}2\mathrm{G}}$)
5:     Project the ocean state from the lat-lon grid into the mesh nodes: ${\mathcal{E}^{\mathrm{G}2\mathrm{M}}_{f}}^{\prime}=\mathrm{ESMLP}(\mathcal{V}_{f}^{G},h,\mathcal{E}^{\mathrm{G}2\mathrm{M}}_{f})$, $h^{\prime}=\mathrm{MLP}_{e1}(h,\sum{\mathcal{E}^{\mathrm{G}2\mathrm{M}}_{f}}^{\prime})$
6:     Update grid node feature: ${\mathcal{V}_{f}^{G}}^{\prime}=\mathrm{MLP}_{e2}(\mathcal{V}_{f}^{G})$
7:     Apply residual connection to update the feature of grid to mesh edge, mesh node, and grid node again: ${\mathcal{E}^{\mathrm{G}2\mathrm{M}}_{f}}={\mathcal{E}^{\mathrm{G}2\mathrm{M}}_{f}}+{\mathcal{E}^{\mathrm{G}2\mathrm{M}}_{f}}^{\prime}$, $h^{\prime}=h+h^{\prime}$, ${\mathcal{V}_{f}^{G}}={\mathcal{V}_{f}^{G}}+{\mathcal{V}_{f}^{G}}^{\prime}$
8:     Multi-scale Interactive Massaging (MIM)
9:     Apply physical process-based edge update module (PEM) to update edge feature: $\mathcal{E}_{f}^{\prime}=PEM(\mathcal{E}_{f},h_{s},h_{r})$
10:     Apply multi-scale node update module (MNM) to update mesh node feature: $h^{\prime}=MNM(h,\sum{\mathcal{E}_{f}}^{\prime})$
11:     Apply residual connection to update the feature of edge and mesh node: $\mathcal{E}_{f}=\mathcal{E}_{f}+\mathcal{E}_{f}^{\prime}$, $h=h+h^{\prime}$
12:     Decoder Project the feature from mesh back to lat-lon grid: ${\mathcal{E}^{\mathrm{M}2\mathrm{G}}_{f}}^{\prime}=\mathrm{ESMLP}(\mathcal{V}_{f}^{G},\mathcal{E}_{f},\mathcal{E}^{\mathrm{M}2\mathrm{G}}_{f})$, ${\mathcal{V}_{f}^{G}}^{\prime}={\mathrm{MLP}_{d1}}(\mathcal{V}_{f}^{G},\sum{\mathcal{E}^{\mathrm{M}2\mathrm{G}}_{f}}^{\prime})$, $\mathcal{V}_{f}^{G}=\mathcal{V}_{f}^{G}+\mathcal{V}_{f}^{G^{\prime}}$, $O_{t+1}=\mathrm{MLP_{d2}}({\mathcal{V}_{f}^{G}})$
13:  until converged
14:  return $MIGNN$
Appendix C Model Training
For the ocean-specific multi-stage framework, we apply Q sub-models to progressively capture subtle changes in ocean variables. We train the first sub-model with M steps finetune. For other sub-models, we train them with N steps finetune. In our experiment, we set Q=2, M=6, and N=10 as an example. Specifically, we train the first sub-model, denoted as $\mathcal{F}^{1}_{1}$ with 200 epochs using 1-step supervision. The learning rate is set to 1e-3. And we use the cosine annealing scheduler to adjust the learning rate until the model converges. After that, we finetune $\mathcal{F}^{1}_{1}$ with 10 epochs using 2-step supervision to get $\mathcal{F}^{2}_{1}$. Then, we finetune $\mathcal{F}^{2}_{1}$ with 10 epochs using 3-step supervision to get $\mathcal{F}^{3}_{1}$. The learning rate is set to 1e-6. And we use the cosine annealing scheduler to adjust the learning rate until the model converges. The model $\mathcal{F}^{4}_{1}$, $\mathcal{F}^{5}_{1}$, and $\mathcal{F}^{M}_{1}$ are finetuned same as $\mathcal{F}^{3}_{1}$. Then we freeze the parameters of $\mathcal{F}^{M}_{1}$ and train the second sub-model $\mathcal{F}^{1}_{2}$ with 200 epochs. The learning rate is set to 1e-6. And we use the cosine annealing scheduler to adjust the learning rate until the model converges. In practice, the model converges after a few epochs, and we stop the training early. The the $\mathcal{F}^{N}_{2}$ can be get step by step with same training produce as $\mathcal{F}^{1}_{2}$. Furthermore, the Q-th model $\mathcal{F}^{N}_{Q}$ can be trained step by step if a higher accuarcy is required. We train all baseline models and our NeuralOM using the same training framework. We will open-source all of our materials—including all training codes, detailed training procedures, testing codes, pre-trained weights, training logs, and more.
Appendix D Additional Results
Additional quantitative results (RMSE and ACC across depth levels 109.73 m, 222.48 m, and 453.94 m) are shown in Figure 5. The RMSE and ACC represent the average performance for 240 ICs, starting from Jan. 1, 2020, and the interval of each IC is 1 day. Additional visual results are presented in Figure 6, Figure 7, Figure 8, Figure 9, Figure 10, and Figure 11, covering 17 important ocean variables across simulation time from 10 to 60 days, with the initial condition at Jun. 1, 2020. Specifically, for variables S, T, and SSH, where variations are subtle relative to their baseline values, anomalies are presented to highlight these minor but significant changes. Conversely, the original data are directly displayed for U and V. For clarity, examples of variable naming conventions are as follows: SSSa denotes sea surface salinity anomaly, Sa109 denotes salinity anomaly at 109.73 m, Sa222 denotes salinity anomaly at 222.48 m, and Sa453 denotes salinity anomaly at 453.94 m. U0 denotes the sea surface stream zonal velocity, U109 denotes the sea stream zonal velocity at 109.73 m, U222 denotes the sea stream zonal velocity at 222.48 m, and U453 denotes the sea stream zonal velocity at 453.94 m. Collectively, these supplementary results further demonstrate the efficacy of the proposed NeuralOM.
Figure 5: The latitude-weighted RMSE (lower is better) and ACC (higher is better) results of several important ocean variables (depth levels at 109.73 m, 222.48 m, and 453.94 m).
Figure 6: 10-day simulation results of different models.
Figure 7: 20-day simulation results of different models.
Figure 8: 30-day simulation results of different models.
Figure 9: 40-day simulation results of different models.
Figure 10: 50-day simulation results of different models.
Figure 11: 60-day simulation results of different models.